# peapods: A Rust-Accelerated Monte Carlo Package for Ising Spin Systems

### v2

@@ 1 Introduction
Rust offers an alternative: compiled performance competitive with C, strong memory safety guarantees enforced at compile time, and ergonomic tooling. The PyO3 library [2] and Maturin build tool [16] provide a mature ecosystem for writing Python extensions in Rust. In this work, we describe peapods, a package that places the entire sampling loop in Rust while exposing a clean Python interface for model construction, analysis, and visualization. The lattice geometry is fully general: users specify neighbor offsets as integer displacement vectors, enabling simulation on hypercubic, triangular, FCC, and other Bravais lattices.
@@ 2 Model
where $\sigma_{i}\in\{-1,+1\}$ are the spin variables, the sum runs over nearest-neighbor pairs, and $J_{ij}$ are the coupling constants. The lattice geometry is defined by a shape (e.g. $L\times L$) and a set of $n$ integer offset vectors specifying the forward neighbor directions; hypercubic lattices use unit vectors along each axis ($n=d$), while the triangular lattice uses offsets $\{(1,0),(0,1),(1,-1)\}$ ($n=3$, coordination $z=6$). The couplings can be specified as an arbitrary user-provided array of shape $(\text{lattice\_shape},n)$, or via one of three built-in distributions:
@@ 3.2 Cluster updates
The Wolff algorithm [5] grows a single cluster from a random seed spin via breadth-first search. Each aligned neighbor is added with probability $p=1-e^{-2J_{ij}\sigma_{i}\sigma_{j}/T}$ when $J_{ij}\sigma_{i}\sigma_{j}>0$. The entire cluster is then flipped deterministically. While each Wolff step flips fewer spins than Swendsen–Wang, it avoids the overhead of a full union-find pass over the lattice.
@@ 3.4 Replica cluster moves
The package implements three replica cluster algorithms for spin glasses, all operating on pairs of replicas at the same temperature. The common framework is:
  1. 1.
    For each temperature, randomly pair the replicas. For each pair $(a,b)$, identify an eligible subgraph based on the overlap $\sigma_{i}^{(a)}\sigma_{i}^{(b)}$.
  2. 2.
    Build clusters on this subgraph (via BFS for Wolff-style single-cluster, or union-find for Swendsen–Wang-style global decomposition).
  3. 3.
    Exchange the spins in selected clusters between the two replicas: $\sigma_{i}^{(a)}\leftrightarrow\sigma_{i}^{(b)}$ for all $i$ in the cluster.
@@
The Houdayer isoenergetic cluster move [7] restricts to the negative-overlap sites $\mathcal{N}=\{i:\sigma_{i}^{(a)}\neq\sigma_{i}^{(b)}\}$ and activates all bonds deterministically (probability 1). This exchange is isoenergetic: because every bond crossing the cluster boundary connects a site inside $\mathcal{N}$ (where the replicas disagree) to a site outside $\mathcal{N}$ (where they agree), the local field at each boundary site is unchanged by the swap. The move is always accepted with no Metropolis test, and the cluster growth is temperature-independent, which is particularly valuable at low temperatures where single-spin-flip moves freeze out.
@@
The Chayes–Machta–Redner (CMR) algorithm [9] generalizes the overlap cluster approach by bonding all same-sign overlap sites (not just negative-overlap), with stochastic bond activation at probability $p=1-e^{-4J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}/T}$ on satisfied bonds. This provides a complementary decorrelation channel to the Houdayer and Jörg moves.
All three modes support both Wolff (single-cluster BFS) and Swendsen–Wang (global union-find) scan styles, selectable via the overlap_cluster_mode parameter.
@@ 3.5 Overlap statistics
can be used for finite-size scaling analysis of the spin glass transition, analogously to the ferromagnetic Binder cumulant. The crossing of $g(T)$ for different system sizes locates the critical temperature $T_{c}$ [18].
@@ 4.2 Lattice geometry
The Lattice struct supports arbitrary Bravais lattice geometries specified by a shape (extents along each axis) and a set of $n$ integer offset vectors defining the forward neighbor directions. For hypercubic lattices, these are the $d$ unit vectors; for the triangular lattice in 2D, they are $\{(1,0),(0,1),(1,-1)\}$. At construction time, the full neighbor table is precomputed: for each of the $N$ sites and each of the $n$ directions, both the forward and backward neighbor flat indices are stored, giving a table of $2nN$ entries (u32). Periodic boundary conditions are applied via rem_euclid on the lattice coordinates.
This precomputed table replaces per-call division and modular arithmetic with a single array lookup, which is faster in practice. The memory cost is modest: for example, 96 KB for a $64\times 64$ triangular lattice ($n=3$) or 6 MB for a $64^{3}$ cubic lattice ($n=3$).
@@ 4.4 Memory management
All buffers (spins, couplings, energies, interactions) are allocated once at initialization and reused across sweeps. The only per-sweep allocations are within cluster updates: union-find arrays and Wolff’s BFS stack, which are allocated per-replica inside the parallel closure. Running statistics are accumulated in f64 to avoid precision loss over many sweeps.
@@ 5 Validation
Figure 1 shows $U_{L}(T)$ for $L=8,16,32,64$ on the square lattice computed with 50,000 sweeps of Metropolis updates combined with Swendsen–Wang cluster updates and parallel tempering. All four curves cross at $T\approx 2.27$ with $U\approx 0.61$, consistent with the exact critical temperature $T_{c}=2/\ln(1+\sqrt{2})\approx 2.269$ [11] and the universal Binder cumulant value of the 2D Ising universality class.
@@
As a test of the general Bravais lattice geometry, we repeat the analysis on the triangular lattice with neighbor offsets $\{(1,0),(0,1),(1,-1)\}$ (coordination $z=6$). The exact critical temperature is $T_{c}=4/\ln 3\approx 3.641$ [12, 13]. For $L=8,16,32$ with 5,000 sweeps, the Binder cumulants cross at $T\approx 3.64$ with $U\approx 0.61$, confirming both the correctness of the neighbor table and the universality of the crossing value (same 2D Ising universality class).
@@ 6 Performance
By moving the entire sampling loop to compiled Rust code, the package avoids the overhead inherent in Python-level Monte Carlo implementations, including per-sweep temporary array allocations, interpreter dispatch in tight loops, and the cost of Python-level iteration over replicas for cluster updates. The Rust backend additionally benefits from on-the-fly neighbor computation (avoiding memory indirection through neighbor tables), Rayon work-stealing parallelism across replicas, and the Xoshiro256** generator which is substantially faster than NumPy’s default RNG for scalar draws. Together, these yield significant speedups over equivalent Python/NumPy implementations, particularly for cluster algorithms where per-temperature Python loops are the bottleneck.
@@ 8 Conclusion
We have presented peapods, a Monte Carlo simulation package for Ising spin systems that combines a Python interface with a Rust computational backend. The package implements the standard toolkit of Metropolis, Gibbs, Swendsen–Wang, Wolff, and parallel tempering algorithms, three replica cluster moves for spin glasses (Houdayer ICM, Jörg, and CMR), overlap statistics for spin glass order parameters, and support for arbitrary Bravais lattice geometries via user-specified neighbor offsets. Replica-level parallelism is achieved via Rayon. The implementation is validated against exact 2D Ising critical points on both the square and triangular lattices. The package is open-source and available on PyPI.

### v3

@@ 1 Introduction
Rust offers an alternative: compiled performance competitive with C, strong memory safety guarantees enforced at compile time, and ergonomic tooling. The PyO3 library [2] and Maturin build tool [16] provide a mature ecosystem for writing Python extensions in Rust. In this work, we describe peapods, a package that places the entire sampling loop in Rust while exposing a clean Python interface for model construction, analysis, and visualization. The lattice geometry is fully general: users specify neighbor offsets as integer displacement vectors, enabling simulation on hypercubic, triangular, FCC, and other Bravais lattices. While the package supports general Ising systems, active development is currently focused on spin glass simulation, where the interplay of parallel tempering, replica overlap cluster moves, and overlap statistics is of particular interest.
Related work
Several existing tools address parts of this space. The ALPS spinmc application [17] provides a mature C++ Monte Carlo framework but does not support spin-glass couplings, requires a heavy build toolchain, and has seen limited maintenance. tamc is a Rust-based Monte Carlo CLI for classical spin systems but provides no cluster algorithms and no Python API. SpinMonteCarlo.jl offers cluster updates in Julia but lacks parallel tempering, spin-glass support, and Python interoperability. Numerous pedagogical Python repositories implement Metropolis-only sampling on square lattices but do not scale to research-grade simulations. To our knowledge, no existing pip-installable package offers a general-purpose Ising Monte Carlo toolkit combining cluster algorithms, parallel tempering, and overlap cluster moves on arbitrary lattice geometries.
@@ 2 Model
where $\sigma_{i}\in\{-1,+1\}$ are the spin variables, the sum runs over nearest-neighbor pairs, and $J_{ij}$ are the coupling constants. The lattice geometry is defined by a shape (e.g. $L\times L$) and a set of $n$ integer offset vectors specifying the forward neighbor directions; hypercubic lattices use unit vectors along each axis ($n=d$), while the triangular lattice uses offsets $\{(1,0),(0,1),(1,-1)\}$ ($n=3$, coordination $z=6$). The FCC lattice uses six forward offsets ($n=6$, coordination $z=12$) and the BCC lattice uses four forward offsets ($n=4$, coordination $z=8$); both are available as named presets via the geometry parameter alongside triangular and hypercubic. The couplings can be specified as an arbitrary user-provided array of shape $(\text{lattice\_shape},n)$, or via one of three built-in distributions:
@@ 3.2 Cluster updates
The Wolff algorithm [5] grows a single cluster from a random seed spin via depth-first search. Each aligned neighbor is added with probability $p=1-e^{-2J_{ij}\sigma_{i}\sigma_{j}/T}$ when $J_{ij}\sigma_{i}\sigma_{j}>0$. The entire cluster is then flipped deterministically. While each Wolff step flips fewer spins than Swendsen–Wang, it avoids the overhead of a full union-find pass over the lattice.
@@ 3.4 Replica cluster moves
The package implements three replica cluster algorithms for spin glasses, all operating on pairs of replicas at the same temperature. For each temperature, replicas are randomly paired; for each pair, clusters are built and flipped. All three modes support both Wolff (single-cluster DFS) and Swendsen–Wang (global union-find) scan styles, and multiple modes can be alternated round-robin via + syntax (e.g. "cmr+houdayer").
@@
The Houdayer isoenergetic cluster move [7] restricts to the negative-overlap sites $\mathcal{N}=\{i:\sigma_{i}^{(a)}\neq\sigma_{i}^{(b)}\}$ and activates all bonds deterministically (probability 1). The move flips both replicas on each selected cluster. This exchange is isoenergetic: because every bond crossing the cluster boundary connects a site inside $\mathcal{N}$ (where the replicas disagree) to a site outside $\mathcal{N}$ (where they agree), the local field at each boundary site is unchanged. The move is always accepted with no Metropolis test, and the cluster growth is temperature-independent, which is particularly valuable at low temperatures where single-spin-flip moves freeze out.
@@
The Chayes–Machta–Redner (CMR) algorithm [9] uses a two-phase construction based on a decomposition of the joint Boltzmann weight of two replicas. Define $r=e^{-2|J_{ij}|/T}$ for each edge. An edge is doubly satisfied if both replicas satisfy it ($J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}>0$ and $J_{ij}\sigma_{i}^{(b)}\sigma_{j}^{(b)}>0$), and singly satisfied if exactly one replica satisfies it.
Phase 1 (blue clusters): Place bonds on doubly-satisfied edges with probability $1-r^{2}$. These define blue clusters. In Swendsen–Wang mode, each non-singleton blue cluster is flipped (both replicas negated) with probability $1/2$. In Wolff mode, the seed’s blue cluster is always flipped.
Phase 2 (grey clusters): Extend the blue cluster decomposition by adding red bonds on singly-satisfied edges (evaluated on post-blue-flip spins) with probability 1r1-r. The grey clusters are the connected components of Blue $\cup$ Red, so blue $\subseteq$ grey. In Swendsen–Wang mode, each non-singleton grey cluster independently flips each replica with probability $1/2$ (i.e. $k\in\{0,1,2,3\}$ uniformly). In Wolff mode, the seed’s grey cluster is flipped with $k\in\{1,2,3\}$ uniformly.
Sites in blue clusters are intentionally flipped in both phases: the composition of the blue and grey flips is the correct update in the CMR graphical representation. The key invariant enabling in-place extension is that blue flips negate both replicas, which swaps which replica is satisfied on a singly-satisfied edge but preserves the singly-satisfied classification itself, so red bonds can be evaluated on post-blue-flip spins without cloning.
@@ 3.5 Overlap statistics
The link overlap provides a complementary bond-level measure of replica similarity:
$q_{l}=\frac{1}{N_{b}}\sum_{\langle i,j\rangle}\bigl(\sigma_{i}^{(a)}\sigma_{j}^{(a)}\bigr)\bigl(\sigma_{i}^{(b)}\sigma_{j}^{(b)}\bigr),$ (5)
where $N_{b}$ is the number of bonds. Unlike the site overlap $q$, the link overlap is invariant under a global spin flip of one replica and is sensitive to the local bond structure of the spin glass state.
@@
can be used for finite-size scaling analysis of the spin glass transition, analogously to the ferromagnetic Binder cumulant. The crossing of $g(T)$ for different system sizes locates the critical temperature $T_{c}$ [19]. The analogous Binder ratio for the link overlap $q_{l}$ is also computed. Additionally, the full overlap histogram $P(q)$ is accumulated per temperature, enabling analysis of the overlap distribution and replica symmetry breaking.
3.6 Cluster size distributions
During Swendsen–Wang and Wolff cluster updates, the package collects Fortuin–Kasteleyn cluster size histograms $n_{s}(T)$, where $n_{s}$ counts the number of clusters of size $s$. Similarly, overlap cluster size histograms are collected during replica cluster moves. From these, the susceptibility-weighted mean cluster size
$\langle s\rangle=\frac{\sum_{s}s^{2}\,n_{s}}{\sum_{s}s\,n_{s}}$ (7)
can be computed, which diverges at the percolation threshold and serves as a diagnostic for the onset of critical behavior.
3.7 Diagnostics
Equilibration diagnostic $\Delta(t)$.
Verifying that a simulation has reached thermal equilibrium is critical for disordered systems where thermalization times are a priori unknown. Following Zhu et al. [20], we compute the equilibration diagnostic
$\Delta(t)=e(t)+\langle J^{2}\rangle\,\beta z\,\bigl(1-q_{l}(t)\bigr),$ (8)
where $e(t)$ is the running average energy per spin, $q_{l}(t)$ is the running average link overlap, $\beta=1/T$, and $z$ is the coordination number. The quantity $\Delta$ vanishes identically in thermal equilibrium, so $\Delta(t)\to 0$ signals that equilibration has been achieved. The diagnostic is evaluated at logarithmically spaced checkpoints ($t=128,256,512,\ldots$) to efficiently monitor convergence over many decades of simulation time.
Integrated autocorrelation time $\tau_{\text{int}}$.
Autocorrelation times quantify how many sweeps are needed to obtain an independent sample. We compute the integrated autocorrelation time via Sokal’s automatic windowing method [21]:
$\tau_{\text{int}}=\frac{1}{2}+\sum_{w=1}^{W}\Gamma(w),$ (9)
where $\Gamma(w)$ is the normalized autocorrelation function and the summation window $W$ is truncated at $w\geq 5\tau_{\text{int}}$ to control noise. The autocorrelation is computed for $m^{2}$ and $q^{2}$ via a streaming ring buffer of $O(\text{max\_lag}\times K)$ memory, where $K$ is the number of temperatures. The autocorrelation time is extracted per realization and then averaged across disorder realizations.
@@ 4.2 Lattice geometry
The Lattice struct supports arbitrary Bravais lattice geometries specified by a shape (extents along each axis) and a set of $n$ integer offset vectors defining the forward neighbor directions. For hypercubic lattices, these are the $d$ unit vectors; for the triangular lattice in 2D, they are $\{(1,0),(0,1),(1,-1)\}$. The FCC lattice ($n=6$, $z=12$) and BCC lattice ($n=4$, $z=8$) are predefined alongside triangular and hypercubic via the geometry parameter. At construction time, the full neighbor table is precomputed: for each of the $N$ sites and each of the $n$ directions, both the forward and backward neighbor flat indices are stored, giving a table of $2nN$ entries (u32). Periodic boundary conditions are applied via rem_euclid on the lattice coordinates.
This precomputed table replaces per-call division and modular arithmetic with a single array lookup, which is faster in practice. The memory cost is modest: for example, 96 KB for a $64\times 64$ triangular lattice ($n=3$), 6 MB for a $64^{3}$ cubic lattice ($n=3$), or 12 MB for a $64^{3}$ FCC lattice ($n=6$).
@@ 4.4 Memory management
All buffers (spins, couplings, energies, interactions) are allocated once at initialization and reused across sweeps. The only per-sweep allocations are within cluster updates: union-find arrays and Wolff’s DFS stack, which are allocated per-replica inside the parallel closure. Running statistics are accumulated in f64 to avoid precision loss over many sweeps.
@@ 5 Validation
Figure 1 shows $U_{L}(T)$ for $L=8,16,32,64$ on the square lattice computed with 500,000 sweeps of Metropolis updates combined with Wolff cluster updates and parallel tempering. All four curves cross at $T\approx 2.27$ with $U\approx 0.61$, consistent with the exact critical temperature $T_{c}=2/\ln(1+\sqrt{2})\approx 2.269$ [11] and the universal Binder cumulant value of the 2D Ising universality class.
@@
As a test of the general Bravais lattice geometry, we repeat the analysis on the triangular lattice with neighbor offsets $\{(1,0),(0,1),(1,-1)\}$ (coordination $z=6$). The exact critical temperature is $T_{c}=4/\ln 3\approx 3.641$ [12, 13]. For $L=8,16,32$ with 500,000 sweeps, the Binder cumulants cross at $T\approx 3.64$ with $U\approx 0.61$, confirming both the correctness of the neighbor table and the universality of the crossing value (same 2D Ising universality class).
@@ 6 Performance
By moving the entire sampling loop to compiled Rust code, the package avoids the overhead inherent in Python-level Monte Carlo implementations, including per-sweep temporary array allocations, interpreter dispatch in tight loops, and the cost of Python-level iteration over replicas for cluster updates. The Rust backend additionally benefits from precomputed neighbor tables (eliminating per-access modular arithmetic), Rayon work-stealing parallelism across replicas, and the Xoshiro256** generator which is substantially faster than NumPy’s default RNG for scalar draws. Together, these yield significant speedups over equivalent Python/NumPy implementations, particularly for cluster algorithms where per-temperature Python loops are the bottleneck.
@@ 7 Usage
# Equilibration diagnostic and autocorrelation time
sg.sample(10000, sweep_mode="metropolis",
pt_interval=1, houdayer_interval=1,
equilibration_diagnostic=True,
autocorrelation_max_lag=200)
sweeps, delta = sg.equilibration_delta()
tau = sg.mags2_tau # integrated autocorrelation time
@@ 8 Conclusion
We have presented peapods, a Monte Carlo simulation package for Ising spin systems that combines a Python interface with a Rust computational backend. The package implements the standard toolkit of Metropolis, Gibbs, Swendsen–Wang, Wolff, and parallel tempering algorithms, three replica cluster moves for spin glasses (Houdayer ICM, Jörg, and CMR with proper two-phase blue/red/grey construction), site and link overlap statistics with full overlap histograms for spin glass order parameters, cluster size distributions, equilibration diagnostics [20], integrated autocorrelation times [21], and support for arbitrary Bravais lattice geometries—including predefined FCC and BCC lattices—via user-specified neighbor offsets. Overlap modes can be alternated round-robin for hybrid update schemes. Replica-level parallelism is achieved via Rayon. The implementation is validated against exact 2D Ising critical points on both the square and triangular lattices. The package is open-source and available on PyPI.